OpenStudy (anonymous):

Help please.. The function f(t) = 3 cos(pi over 6t) + 5 represents the tide in Blastic Sea. It has a maximum of 8 feet when time (t) is 0 and a minimum of 2 feet. The sea repeats this cycle every 12 hours. After nine hours, how high is the tide? 12 feet 5 feet 4.5 feet 2.5 feet

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@jim_thompson5910

zepdrix (zepdrix):

So they just want us to evaluate this function at t=9 hours.\[\Large\rm f(\color{orangered}{t})=3\cos\left(\frac{\pi}{6}\color{orangered}{t}\right)+5\]\[\Large\rm f(\color{orangered}{9})=3\cos\left(\frac{\pi}{6}\color{orangered}{9}\right)+5\]

zepdrix (zepdrix):

Looks like the inside simplifies a lil bit,\[\Large\rm f(9)=3\cos\left(\frac{9\pi}{6}\right)+5\]\[\Large\rm f(9)=3\cos\left(\frac{3\pi}{2}\right)+5\]Do you remember your unit circle? :O cos(3pi/2) = ?

OpenStudy (anonymous):

I do not.. I can put it into Mathway if you like :oo

zepdrix (zepdrix):

You... -_- Yah do that then :p throw it into a calculator, you should get a nice simple answer.

OpenStudy (anonymous):

0 correct?

zepdrix (zepdrix):

yes. good.

OpenStudy (anonymous):

Whoop ^.^

zepdrix (zepdrix):

\[\Large\rm f(9)=3\cdot0+5\]

OpenStudy (anonymous):

so f(9)=5

zepdrix (zepdrix):

yay good job \c:/

OpenStudy (anonymous):

And that's my answer ^.^ Thank you!!